Factor completely Please show work as I am not sure how to do this.
9t2 + 14t + 5

Answer by Alan3354(69443)

Factor completely Please show work as I am not sure how to do this.
9t2 + 14t + 5
-------------
It will be (at + b)*(ct + d)
b*d = 5
--> (at + 1)*(bt + 5)
Find integers for a & b so that a*b = 9 and 5a + b = 15
It's trial and error with only a few possibilities.
-----
For a & b, it's either 1*9 or 3*3
==========
There's nothing to be gained by me doing it.